OpenStudy (anonymous):

A shirt that normally sells for $28 is on sale at a 10% discount. What is the sale price of the shirt?

OpenStudy (anonymous):

$25.20

OpenStudy (anonymous):

28- 10 % of 28=?

OpenStudy (anonymous):

yes but I do it a little different

OpenStudy (anonymous):

or 90% of 28

OpenStudy (anonymous):

$28 x 10% = 2.8 2.8- 28$ = 25.20

OpenStudy (anonymous):

10% of 28 lol I used to think that too but no

OpenStudy (anonymous):

28-2.80=25.20